OpenStudy (anonymous):

Solve the absolute value equation and check: |6x-4|-8=12

hartnn (hartnn):

first add 8 to both sides, what u get ?

hartnn (hartnn):

did u get |6x-4| = 20 ??

OpenStudy (anonymous):

Yes I got 20 but now I'm trying to figure out the other part...this is where I get confused.

hartnn (hartnn):

ok, so if |a|=b then a=b OR a=-b here |6x-4| = 20 so 6x-4 = 20 OR 6x-4=-20 can u solve these 2 equations ??

OpenStudy (anonymous):

hmm I'm still confused, math is not my subject ,I do much better with english ;) I'm thinking I need two solutions solving the absolute first...

hartnn (hartnn):

yes, u will get 2 solution one when solving for x in 6x-4 = 20 and other when solving for x in 6x-4 = -20 can u solve 6x-4 = 20 ?? hint :1st add 4 to both sides.

OpenStudy (anonymous):

So I end up with -24/6 with x=4 for one, is this correct?

hartnn (hartnn):

nopes, for 6x-4 = 20 u get 6x= 20+4= 24 x=24/6 = 4 ok ?

OpenStudy (anonymous):

Ok and then how do I check the answer?

hartnn (hartnn):

wait, 2nd equation is still remaining. 6x-4=-20 add 4 6x = -20+4 6x = -16 x= (-16/6) = -8/3 ok with this ?

hartnn (hartnn):

please ask if u do not follow any step.

OpenStudy (anonymous):

This equation I'm having a hard time understanding. so how did you get -16 and the -8/3?

OpenStudy (anonymous):

oh I think I just figured it out except the -8/3 ;)

hartnn (hartnn):

-20+4 =-16 ok, 20 -4 = 16 right ? so -(20-4) = -16 -20+4 = -16 ok with this ? and 16/6 = (8*2)/(3*2) = 8/3 2 getting cancelled from numerator and denominator.

OpenStudy (anonymous):

ok so we just round up to the 3?

OpenStudy (anonymous):

I get it now that I'm writing it down seeing the steps of how you ended up with each solution. Our instructor did not explain each step...it looks like a couple of the steps were skipped. I need to know each step since I'm not good at this. Thank you so much!

hartnn (hartnn):

welcome :) write and ask if there's still any doubts in any step...

hartnn (hartnn):

still checking is remaining \(\checkmark \) take one value which u got and put it again in left side of equation and see whether u get the right side. so, lets take -8/3 so |6x-4|-8 =|6(-8/3)-4|-8 =|-16-4|-8 =|-20|-8 =20-8 because |-a|=a =12 so u get right side so (-8/3) is \(\checkmark \) can u do the same for x=4 now ??

hartnn (hartnn):

@hills i assume u are trying.

OpenStudy (anonymous):

lol Yes l'm just about done and then I will compare with yours. I was helping someone on another subject ;)

hartnn (hartnn):

thats great :)

OpenStudy (anonymous):

|6(4)-4|-8 =|24-4|-8 =|20|-8 =20-8 =12 so 4 is correct, right?

hartnn (hartnn):

yes, u got it :) \(\huge\checkmark\)

OpenStudy (anonymous):

Yay! It sure helped following step by step ;) Again, thank you so much for your help!

hartnn (hartnn):

welcome :)